Please help me solve this problem.
C= 3x + 2y

Constraints:
x + y <= 6
-x + y >= 0
x >= 0

Answer by stanbon(48545)

Graph the line x+y=6
Graph the line y=x
Graph the line x=0
Shade the area enclosed by the three lines and find the point
where y=x intersects the line x+y=6
The three vertices of the triangle formed by the shaded area are
(0,0), (0,6), and (3,3)
Evaluate "C" at each of these pairs of values.
For example, using the pair (3,3) the value of
the cost would be as follows:
Cost = 3x+2y
Cost = 3(3)+2(3) = 15
Do the same thing with (0,0) and (0,6)
Cheers,
stan H.